# Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions

### v2

@@ Abstract
We study the bidding problem in repeated uniform price multi-unit auctions from the perspective of a single value-maximizing buyer who aims to maximize their cumulative value over $T$ rounds while adhering to return-on-investment (RoI) constraints in each round. Buyers adopt $m$-uniform bidding format, where they submit $m$ bid-quantity pairs $(b_{i},q_{i})$ to demand $q_{i}$ units at bid $b_{i}$. We introduce safe bidding strategies as those that satisfy RoI constraints in every auction, regardless of competing bids. We show that these strategies depend only on the bidder’s valuation curve, and the bidder can focus on a finite subset of this class without loss of generality. While the number of strategies in this subset is exponential in $m$, we develop a polynomial-time algorithm to learn the optimal safe strategy that achieves sublinear regret in the online setting, where regret is measured against a clairvoyant benchmark that knows the competing bids a priori and selects a fixed hindsight optimal safe strategy. We then evaluate the performance of safe strategies against a clairvoyant that selects the optimal strategy from a richer class of strategies in the online setting. In this scenario, we compute the richness ratio, $\alpha\in(0,1]$ for the class of strategies chosen by the clairvoyant and show that our algorithm, designed to learn safe strategies, achieves $\alpha$-approximate sublinear regret against these stronger benchmarks. Experiments on semi-synthetic data from real-world auctions show that safe strategies substantially outperform the derived theoretical bounds, making them quite appealing in practice.
@@ 1.  Introduction
In a uniform price multi-unit auction, the auctioneer sells $K$ identical units of a single good to buyers who may demand multiple units, with the per-unit price set at the $K^{th}$ highest bid. These auctions are widely used to allocate scarce resources in critical markets such as emissions permits, energy markets, and Treasury auctions. To succeed in these markets, bidders must develop effective bidding strategies that balance long-term value maximization with financial risk management, all while operating under limited information, uncertainty in competing bids and strategic (or adversarial) behavior of other bidders, among other challenges.
In this paper, we model the bidders as a value-maximizing agent with per-round return-on-investment (RoI) constraints. This model reflects real-world decision-making in industries where managing financial risks is as critical as maximizing value. Unlike traditional mechanism design, which assumes bidders with quasilinear utility (i.e., utility decreases linearly with payments), many practical settings involve bidders who optimize total value while adhering to financial constraints such as RoI or budgets. These scenarios frequently arise in industries where agents or algorithms bid on behalf of clients, optimizing for high-level objectives within strict constraints—a context akin to the principal-agent framework (Fadaei and Bichler, 2016; Aggarwal et al., 2024). Examples include autobidders in online advertising, where advertisers seek to maximize clicks while keeping the average cost per click below a threshold (Lucier et al., 2024; Balseiro et al., 2021a; Deng et al., 2023b), and consultants bidding for small firms in emission permit auctions (EEX, 2024).
Building on this, we study the bidding problem in repeated uniform price auctions over $T$ rounds, from the perspective of a single bidder who seeks to maximize cumulative value while adhering to per-round RoI constraints—that is, ensuring that the value obtained in each round is at least a fixed multiple of the corresponding payment (see Eq. 2).
To address practical bidding interfaces, we assume bidders adopt $m$-uniform bidding format, a generalization of the uniform bidding format (De Keijzer et al., 2013; Birmpas et al., 2019) for some $m\in\mathbb{N}$. In a $m$-uniform strategy $\mathbf{b}:=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$, bidders bid $b_{1}$ for the first $q_{1}$ units, $b_{2}$ for the next $q_{2}$ units and so on (see Definition 1).
@@ 1.1.  Our Contributions
Safe Bidding Strategies (Section 3). To ensure that the bidder satisfies the RoI constraint without knowing the competing bids, we introduce the concept of safe bidding strategies. These strategies guarantee that the RoI constraint is met regardless of how the other participants bid. We then identify the class of $m$-uniform safe undominated bidding strategies, denoted by $\mathscr{S}_{m}^{\star}$ (Theorem 3.2), and demonstrate that the strategies within this class depend solely on the bidder’s valuation vector $[v_{1},\dots,v_{K}]$ and exhibit a “nested” structure (illustrated in Fig. 1).
@@
Figure 1: Nested structure of the bidding strategies in $\mathscr{S}_{3}^{\star}$. Consider the strategy $\mathbf{b}=\langle(w_{3},3),(w_{7},4),(w_{9},2)\rangle\in\mathscr{S}_{3}^{\star}$, where we note that $Q_{1}=3$, $Q_{2}=3+4=7$ and $Q_{3}=3+4+2=9$. The $j^{th}$ highest bid (i.e., $b_{j}$) is the average of the first $Q_{j}$ entries of the valuation vector, i.e., $b_{j}=w_{Q_{j}}$, where $w_{j}=\frac{1}{j}\sum_{\ell\leq j}v_{\ell}$.
@@
Learning Safe Bidding Strategies (Section 4). Designing an algorithm to learn the optimal $m$-uniform safe bidding strategy poses significant challenges due to the exponential size of the decision space in $m$. To tackle this, we first consider an offline setting where the competing bids are known in advance. In this setting, we construct a polynomial-sized directed acyclic graph (DAG) with carefully assigned edge weights and show that determining the maximum weight path in the DAG is equivalent to computing the optimal $m$-uniform safe bidding strategy (Theorem 4.2).
Building on this, we develop an online learning algorithm when the competing bids are unknown. [^1] For a detailed discussion on leveraging offline algorithms for no-regret learning, we refer readers to Roughgarden and Wang (2019); Niazadeh et al. (2022); Brânzei et al. (2023). Our algorithm runs in polynomial time and achieves a regret of $\widetilde{O}(M\sqrt{mT})$ in a full-information setting and $\widetilde{O}(M^{3/2}m^{3/2}\sqrt{T})$ regret in a bandit setting, where the clairvoyant selects the fixed hindsight optimal safe strategy (Theorem 4.3). Additionally, we establish a regret lower bound of $O(M\sqrt{T})$ (Theorem 4.4).
@@
The problem of learning to bid in multi-unit uniform price and pay-as-bid auctions has been explored recently by Brânzei et al. (2023); Galgana and Golrezaei (2024); Potfer et al. (2024), assuming the bidders are quasilinear utility maximizers. Our work differs from them in two key ways: (a) we consider value-maximizing buyers with RoI constraints, a fundamentally different behavioral model, and (b) from a technical perspective, unlike prior works that require bid spaces to be discretized, our approach does not, due to the structure of safe strategies. Consequently, the time complexity becomes independent of $T$ in both online and offline settings  [^2] Assuming the discretization level in the offline setting is same as that in the online setting. In the online setting, time complexity refers to per-round running time., and under bandit feedback, the dependence on $T$ in the regret bound improves from $T^{2/3}$ to $\sqrt{T}$ implying that it is easier to learn safe strategies for value maximizers compared to bidding strategies for quasilinear utility maximizers.
Richer Classes of Strategies for the Clairvoyant (Section 5). A key novel aspect of our work is to evaluate the performance of safe bidding strategies with at most $m$ bid-quantity pairs and the online learning algorithm from Section 4 against a clairvoyant that selects the optimal strategy from a richer class of strategies. In this case, the bidder (learner) achieves $\alpha$-approximate sublinear regret (same as Theorem 4.3) for a parameter $\alpha\in(0,1]$—defined as the richness ratio—thereby demonstrating the robustness of the safe bidding strategies against a broad range of benchmarks (Corollary 5.2). One of our main contributions is to compute the richness ratio, $\alpha$, for various classes of bidding strategies. Specifically, when the clairvoyant selects the optimal bidding from the class of
@@
(a) strategies that are RoI-feasible, not necessarily safe and have at most $m$ bid-quantity pairs, we get $\alpha=\frac{1}{2}$, notably independent of $m$ (Theorem 5.3).
@@
(b) safe strategies with at most $m^{\prime}$ bid-quantity pairs, where $m^{\prime}\geq m$, we obtain $\alpha=\frac{m}{m^{\prime}}$ (Theorem 5.4).
@@
(c) strategies that are RoI-feasible, not necessarily safe and have at most $m^{\prime}$ bid-quantity pairs, where $m^{\prime}\geq m$, we get that $\alpha=\frac{m}{2m^{\prime}}$ (Theorem 5.5).
@@
Computing $\alpha$ involves two key challenges: (a) deriving an upper bound on the ratio between the value obtained by the clairvoyant’s optimal strategy and that of the optimal safe strategy in the worst case (see Definition 4 for details), and (b) proving that this upper bound is tight. The key ideas for deriving the upper bound are outlined in Section 5. The construction of problem instances, each tailored to a specific strategy class, is non-trivial due to their exponential size in $m$ and the need for a careful choice of competing bids and the valuation vector, as detailed in Sections D.1 to D.4. Simulations using semi-synthetic data show that the empirical richness ratios in practical scenarios are significantly better than the theoretical bounds, making safe strategies highly appealing in practice (Section 5.3).
@@ 1.2.  Related Work
Value Maximizers and RoI Constraints. The concept of agents as value maximizers within financial constraints is a well-established notion in microeconomic theory (Mas-Colell et al., 1995, pp. 50). In mechanism design literature, one of the earliest explorations of value-maximizing agents was conducted by Wilkens et al. (2016). Their work primarily delved into the single-parameter setting, characterizing truthful auctions for value maximizers. Similarly, Fadaei and Bichler (2016) and Lu et al. (2023) studied truthful (approximate) revenue-maximizing mechanisms in combinatorial markets tailored for such agents.
@@
In recent years, there has been growing interest in RoI-constrained value maximizers, particularly in the context of autobidding and online advertising. One of the earliest works in this area is Golrezaei et al. (2021), which studied auction design for RoI-constrained buyers and validated the presence of such soft financial constraints using data from online ad auctions. Broadly, the literature in this space can be divided into two categories: (i) works that design optimal auctions under RoI constraints (Balseiro et al., 2021a, b, 2022; Deng et al., 2021, 2023b), and (ii) works that characterize optimal bidding strategies and/or develop learning algorithms in repeated auction settings (Aggarwal et al., 2019; Deng et al., 2023a; Golrezaei et al., 2023; Castiglioni et al., 2024; Aggarwal et al., 2025; Lucier et al., 2024). Our work primarily aligns with the latter. We study value-maximizing buyers in uniform price auctions under RoI constraints and demonstrate that these buyers can employ safe bidding strategies that are efficiently learnable and robust against a variety of strong benchmarks.
@@
Multi-unit Auctions. In this work, we focus on a subset of combinatorial auctions termed as multi-unit auctions in which multiple identical goods are sold to a group of buyers. [^3] For a comprehensive survey on combinatorial auctions, we refer the readers to several excellent works by De Vries and Vohra (2003); Blumrosen and Nisan (2007); Palacios-Huerta et al. (2022). These auctions find widespread application in various practical scenarios, including Treasury auctions (Hortaçsu and McAdams, 2010), procurement auctions (Cramton and Ausubel, 2006), electricity markets (Tierney et al., 2008), and emissions permit auctions (Schmalensee and Stavins, 2017). While several works have focused on studying the equlibria properties in these auctions (Ausubel et al., 2014; Markakis and Telelis, 2015; De Keijzer et al., 2013), computing equilibrium strategies is usually intractable in general multi-unit auctions due to multi-dimensional valuations. As a result, a growing body of work has emphasized on designing optimal bidding strategies in a prior-free setting, i.e., without any assumptions of the competing bidders’ bids or valuations (Galgana and Golrezaei, 2024; Potfer et al., 2024; Brânzei et al., 2023). Our work contributes to this literature by learning optimal strategies for value-maximizing buyers in repeated uniform price auctions.

### v4

# Negin Golrezaei

@@ Abstract
We study the bidding problem in repeated uniform price multi-unit auctions, a format prevalent in emission permit auctions, Treasury auctions and energy markets, from the perspective of a value-maximizing buyer. The buyer aims to maximize their cumulative value over $T$ rounds while adhering to per-round return-on-investment (RoI) constraints in a strategic (or adversarial) environment. The buyer uses $m$-uniform bidding format, submitting $m$ bid-quantity pairs $(b_{i},q_{i})$ to demand $q_{i}$ units at bid $b_{i}$, with $m\ll M$ in practice, where, $M$ is the maximum demand of the buyer.
We introduce the notion of safe bidding strategies as those that satisfy the RoI constraint irrespective of competing bids. Despite the stringent requirement, we show that these strategies satisfy a mild no-overbidding condition, depend only on the bidder’s valuation curve, and the bidder can focus on a finite subset without loss of generality. While the number of strategies in this subset is $O(M^{m})$, leveraging a key insight on the decomposition of a strategy’s obtained value across its constituent bid-quantity pairs, we develop a polynomial-time algorithm to learn the optimal safe strategy that achieves a regret of $\widetilde{O}(M\sqrt{mT})$ and $\widetilde{O}(M^{2}m^{3/2}\sqrt{T})$ in the full-information and bandit settings respectively, where regret is measured against a clairvoyant that selects the fixed hindsight optimal safe strategy. We complement the upper bounds by establishing a $O(M\sqrt{T})$ regret lower bound.
We then assess the robustness of safe strategies against the hindsight-optimal strategy from a richer class of strategies. We define the richness ratio $\alpha\in(0,1]$ as the minimum ratio of the value of the optimal safe strategy to that of the optimal strategy from richer class and construct hard instances showing the tightness of $\alpha$. Our algorithm achieves $\alpha$-approximate sublinear regret against these stronger benchmarks. Simulations on semi-synthetic real-world auction data demonstrate that empirical richness ratios are significantly better than the tight theoretical bounds. Finally, we show that (variants of) safe strategies and the proposed learning algorithm are well-suited to cater to more nuanced behavioral models of the buyer and the competitors. The simplicity, robustness, and adaptability of safe bidding strategies make them highly promising for practitioners.
@@ 1. Introduction
In a uniform price multi-unit auction, the auctioneer sells $K$ identical units of a single good to buyers who may demand multiple units, with the per-unit price set at the $K^{th}$ highest bid. These auctions are widely used to allocate scarce resources in critical avenues such as emissions permit auctions, wholesale electricity markets, and Treasury auctions. To succeed in these markets, bidders must develop effective bidding strategies that balance long-term value maximization with financial risk management, all while operating under limited information, uncertainty in competing bids and strategic (or adversarial) behavior of other bidders, among other challenges.
In this paper, we model the bidders as a value-maximizing agent with per-round return-on-investment (RoI) constraints. This model reflects real-world decision-making in industries where managing financial risks is as important as maximizing value. Unlike traditional auction theory, which assumes bidders with quasilinear utility (i.e., utility decreases linearly with payments), many practical settings involve bidders who optimize total value while adhering to financial constraints such as RoI or budgets. These scenarios frequently arise in industries where agents or algorithms bid on behalf of clients, optimizing for high-level objectives within strict constraints–a context akin to the principal-agent framework (Fadaei and Bichler, 2016; Aggarwal et al., 2024).
For example, in online advertising markets, performance marketing agencies often manage autobidding systems for clients who seek to maximize conversions or clicks while ensuring the average cost-per-click or return-on-ad-spend remains within specified limits (Lucier et al., 2024; Balseiro et al., 2021a; Deng et al., 2023b). Similarly, in the EU Emissions Trading System (EU ETS), many compliant companies participate in emission permits auctions indirectly through third-party firms such as MK Brokers JSC [^2] https://mkb.bg/en/services/sustainability-products/., Act Group [^3] https://www.actgroup.com/products-solutions/compliance/emissions-compliance-products., and AFS Group [^4] https://www.afsgroup.nl/exchange-traded-energy-instruments.. These intermediaries bid on behalf of clients often under contractual constraints on allowable expenditures or minimum cost-effectiveness of permit acquisition (Ellerman et al., 2010).
Building on this, we study the bidding problem in repeated uniform price auctions over $T$ rounds, from the perspective of a single bidder who seeks to maximize cumulative value while adhering to per-round RoI constraints—that is, ensuring that the value obtained in each round is at least a fixed multiple of the corresponding payment (see Eq. 2). To address practical bidding interfaces, we assume bidders adopt $m$-uniform bidding format, a generalization of the uniform bidding format (De Keijzer et al., 2013; Birmpas et al., 2019) for some $m\in\mathbb{N}$. In a $m$-uniform strategy $\mathbf{b}:=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$, bidders bid $b_{1}$ for the first $q_{1}$ units, $b_{2}$ for the next $q_{2}$ units and so on (see Definition 1).
@@ 1.1. Our Contributions
Safe Bidding Strategies (Section 3). To ensure that the bidder satisfies the RoI constraint without knowing the (potentially adversarial) competing bids in the online setting, we introduce the concept of safe bidding strategies. These strategies guarantee that the RoI constraint is met regardless of how the other participants bid. We show that these strategies follow a mild no-overbidding condition (see Definition 3) and the bidder can focus on a finite subset of this class without loss of generality. For any $m\in\mathbb{N}$, we characterize this finite subset as the class of $m$-uniform safe undominated bidding strategies, denoted by $\mathscr{S}_{m}^{\star}$ (Theorem 3.2), and demonstrate that the strategies within this class depend solely on the bidder’s valuation vector $v_{1},\dots,v_{K}]$ and exhibit a “nested” structure as illustrated in Fig. 1.
@@
Figure 1: Nested structure of the bidding strategies in $\mathscr{S}_{3}^{\star}$. Consider the strategy $\mathbf{b}=\langle(w_{3},3),(w_{7},4),(w_{9},2)\rangle\in\mathscr{S}_{3}^{\star}$, where we note that $Q_{1}=3$, $Q_{2}=3+4=7$ and $Q_{3}=3+4+2=9$. The $j^{th}$ highest bid (i.e., $b_{j}$) is the average of the first $Q_{j}=\sum_{\ell\leq j}q_{\ell}$ entries of the valuation vector, i.e., $b_{j}=w_{Q_{j}}$, where $w_{j}=\frac{1}{j}\sum_{\ell\leq j}v_{\ell}$.
@@
Learning Safe Bidding Strategies (Section 4). Designing an algorithm to learn the optimal safe bidding strategy with at most $m$ bid-quantity pairs poses significant challenges as the size of the decision space in $O(M^{m})$, where $M$ is the maximum number of units demanded by the bidder and in practice, $m\ll M$. To this end, we first consider an offline setting where the competing bids are known a priori. We show that the value obtained by a $m$-uniform safe strategy can be decomposed across its $m$ constituent bid-quantity pairs. Leveraging this decomposition, we construct a directed acyclic graph (DAG), of size $\text{poly}(m,M)$, with carefully assigned edge weights and show that determining the maximum weight path in the DAG is equivalent to computing the optimal safe bidding strategy with at most $m$ bid-quantity pairs (Theorem 4.2). This reduction allows us to compute the offline optimal safe strategy in $\text{poly}(m,M)$ time, despite the exponential size of the original decision space.
[…]
@@
The problem of learning to bid in multi-unit uniform price and pay-as-bid auctions has been explored recently by Brânzei et al. (2023); Galgana and Golrezaei (2024); Potfer et al. (2024), assuming the bidders are quasilinear utility maximizers. Our work differs from them in two key ways: (a) we consider value-maximizing buyers with RoI constraints, a fundamentally different behavioral model, and (b) from a technical perspective, unlike prior works that require bid spaces to be discretized, our approach does not, due to the structure of safe strategies. Consequently, the time complexity becomes independent of $T$ in both online and offline settings [^7] Assuming the discretization level in the offline setting is the same as that in the online setting. In the online setting, time complexity refers to per-round running time., and under bandit feedback, the dependence on $T$ in the regret bound improves from $T^{2/3}$ to $\sqrt{T}$ implying that it is easier to learn safe strategies for value maximizers compared to bidding strategies for quasilinear utility maximizers.
Richer Classes of Strategies for the Clairvoyant (Section 5). We then evaluate the performance of the bidder’s strategy class, i.e., safe bidding strategies with at most $m$ bid-quantity pairs and the online learning algorithm from Section 4 against a clairvoyant benchmark that selects the optimal strategy from a richer class. Here, a richer class refers to any class of strategies that is a superset of the bidder’s strategy class (see examples below). Recall that the bidder chooses to follow the $m$-uniform safe strategies to satisfy the RoI constraints with limited knowledge about the competing bids in any auction in the online setting.
Comparing against such stronger benchmarks demonstrates the robustness of the bidder’s strategy class. We quantify this robustness via the richness ratio $\alpha\in(0,1]$ (Definition 4), analogous to an approximation ratio. We prove that for any richer class with richness ratio $\alpha$, the bidder achieves $\alpha$-approximate regret (as defined in Eq. 8) of the same order as in Theorem 4.3 (see Corollary 5.2). One of our main contributions is to compute the richness ratio $\alpha$ for various classes of bidding strategies. We define richness along two dimensions: (i) the safety of strategies (safe vs. only RoI-feasible), and (ii) the number of bid-quantity pairs allowed. A strategy is defined as RoI-feasible (or feasible in short) if it satisfies the RoI constraint for a given sequence of competing bids over $T$ rounds, unlike safe strategies that satisfy the constraint for every such sequence. We choose these two dimensions of richness as the bidder follows the class of strategies that are (i) safe and (ii) have at most $m$ bid-quantity pairs.
When the clairvoyant selects the optimal strategy from the class of
  • (a)
@@
    strategies that are RoI-feasible and have at most $m$ bid-quantity pairs, we get $\alpha=\frac{1}{2}$ (Theorem 5.3). Notably, for this particular choice of richer class of strategies, $\alpha$ is independent of $m$.
  • (b)
@@
    safe strategies with at most $m^{\prime}$ bid-quantity pairs, where $m^{\prime}\geq m$, we obtain $\alpha=\frac{m}{m^{\prime}}$ (Theorem 5.4). Here, setting $m^{\prime}=m$ refers to the case considered in Section 4 where the bidder and clairvoyant follow the same class of strategies. In this case, Theorem 5.4 correctly implies the optimal value of $\alpha=1$.
  • (c)
@@
    strategies that are RoI-feasible and have at most $m^{\prime}$ bid-quantity pairs, where $m^{\prime}\geq m$, we get $\alpha=\frac{m}{2m^{\prime}}$ (Theorem 5.5). Setting $m^{\prime}=m$ in this case recovers the result of Theorem 5.3.
@@
Computing the richness ratio involves two different challenges: (a) deriving an upper bound on the ratio between the value obtained by the clairvoyant’s optimal strategy from the richer class and that of the optimal safe strategy in the worst case (see Definition 4 for details), and (b) proving that this upper bound is tight by constructing a hard problem instance for the given richer class of strategies. The key ideas for deriving the upper bounds are outlined in Section 5. The hard problem instances that show the tightness of the bounds are highly non-trivial where we need to consider an instance of size that is exponential in $m$, with careful choice of competing bids and the valuation vector (see details in Sections D.1 to D.3).
Extensions (Section 6). Finally, we show that safe strategies and the learning algorithm from Section 4 can be adapted to handle more nuanced modeling assumptions. First, as competitors may adapt to the bidder’s past actions in repeated auctions, we consider an adaptive adversary model for generating competing bids (as opposed to an oblivious one). In this setting, the learning algorithm with appropriate modifications still yields a high-probability regret bound of $\text{poly}(m,M)\cdot\sqrt{T}$ (Section 6.1). Second, when the RoI constraint is enforced over a sliding window of $T_{0}\in\mathbb{N}$ rounds instead of each round, we propose a heuristic based on shifted safe strategies for this setting that shows good empirical performance on semi-synthetic real-world auction data (Section 6.2). Third, in the case of time-varying valuation vectors, that are either sampled from a distribution or generated adversarially from a finite set $\mathcal{V}$, we treat the valuation vector as contexts and show that a contextual version of our learning algorithms achieves $\text{poly}(m,M,|\mathcal{V}|)\cdot\sqrt{T}$ regret in this setting (Section 6.3). These extensions highlight the practical appeal and flexibility of our approach.
1.2. Managerial Insights
Our findings offer several key implications for RoI-constrained value maximizing buyers.
  1. 1.
    Safe strategies are an intuitive class of bidding strategies due to their mild no-overbidding conditions. As the bidder can focus on a finite subset of the class without any loss of generality, designing and implementing the proposed learning algorithm in repeated auction environments is quite efficient from a practical perspective. Furthermore, the theoretical regret upper and lower bounds with $\sqrt{T}$ dependence indicate that the proposed algorithm is near-optimal, which is desirable to practitioners.
  2. 2.
    We crucially show the robustness of safe strategies and the learning algorithm by proving that they are approximately optimal against stronger benchmarks. A key highlight here is that for any given $m$, the cost of following safe strategies compared to the instance dependent RoI-feasible strategies, which are impossible to know a priori, is uniformly bounded. So, a clairvoyant with prior knowledge of the adversarially generated competing bids in this setting can not be arbitrarily better than the bidder following safe strategies.
  3. 3.
    Simulations on semi-synthetic data from EU Emission Trading System (ETS) auctions show that the empirical richness ratios in real-world scenarios are significantly better than tight theoretical bounds, which reinforces the fact that tight bounds are attained under highly pathological settings and suggests that safe strategies can perform near-optimal in practice, even for small values $m$ (Section 5.3).
  4. 4.
    Finally, we show that variants of safe strategies and the proposed learning algorithm can be appropriately deployed in several practical extensions of the primary model considered in this work. This aspect is particularly interesting as it allows practitioners to tweak safe strategies and/or the proposed algorithm to design novel algorithms or heuristics to adapt to their objectives and risk appetite that may change over time.
@@ 1.3. Related Work
Value Maximizers and RoI Constraints. The concept of agents as value maximizers within financial constraints is a well-established notion in microeconomic theory (Mas-Colell et al., 1995). In mechanism design literature, one of the earliest explorations of value-maximizing agents was conducted by Wilkens et al. (2016). Their work primarily delved into the single-parameter setting, characterizing truthful auctions for value maximizers. Similarly, Fadaei and Bichler (2016) and Lu et al. (2023) studied truthful (approximate) revenue-maximizing mechanisms in combinatorial markets tailored for such agents. More recently, Tang et al. (2024) investigated auction design in a setting where buyers may be either quasilinear utility maximizers or RoI-constrained value maximizers. Xu et al. (2024) proposed the allowance utility maximization model, which generalizes both quasilinear and value-maximizing preferences.
@@
In recent years, there has been growing interest in RoI-constrained value maximizers, particularly in the context of autobidding and online advertising. One of the earliest works in this area is Golrezaei et al. (2021c), which studied auction design for RoI-constrained buyers and validated the presence of such soft financial constraints using data from online ad auctions. Broadly, the literature in this space can be divided into two categories: (i) works that design optimal auctions under RoI constraints (Balseiro et al., 2021a, b, 2022; Deng et al., 2021, 2023b), and (ii) works that characterize optimal bidding strategies and/or develop learning algorithms in repeated auction settings (Aggarwal et al., 2019; Deng et al., 2023a; Golrezaei et al., 2023; Castiglioni et al., 2024; Aggarwal et al., 2025; Lucier et al., 2024). Our work primarily aligns with the latter. We study value-maximizing buyers in uniform price auctions under RoI constraints and show that they can employ safe strategies that are efficiently learnable and robust against a variety of strong benchmarks.
@@
Multi-unit Auctions. In this work, we focus on a subset of combinatorial auctions termed as multi-unit auctions in which multiple identical goods are sold to a group of buyers. [^8] For a comprehensive survey on combinatorial auctions, we refer readers to several excellent works by De Vries and Vohra (2003); Blumrosen and Nisan (2007); Palacios-Huerta et al. (2024). These auctions find widespread application in various practical scenarios, including Treasury auctions (Nyborg et al., 2002; Hortaçsu and McAdams, 2010; Hortaçsu and Kastl, 2012), procurement auctions (Cramton and Ausubel, 2006), electricity markets (Tierney et al., 2008; Fabra et al., 2006), and emissions permit auctions (Goulder and Schein, 2013; Schmalensee and Stavins, 2017; Goldner et al., 2020).
While several works have focused on studying the equilibria properties in these auctions (Ausubel et al., 2014; Engelbrecht-Wiggans and Kahn, 1998; Noussair, 1995; Markakis and Telelis, 2015; De Keijzer et al., 2013), computing equilibrium strategies is usually intractable in general multi-unit auctions due to multi-dimensional valuations (Kasberger and Woodward, 2025). As a result, a growing body of work has emphasized on designing optimal bidding strategies in a prior-free setting, i.e., without any assumptions of the competing bidders’ bids or valuations–in a repeated setting (Galgana and Golrezaei, 2024; Potfer et al., 2024; Brânzei et al., 2023) or under a minimax loss framework (Kasberger and Woodward, 2025). Our work contributes to this literature by learning optimal strategies for value-maximizing buyers in repeated uniform price auctions.
Data-Driven Methods in Auctions. Traditionally, the problems of bidding in auctions and determining parameters such as reserve prices have been studied in the Bayesian framework (Myerson, 1981; Riley and Samuelson, 1981; Hartline and Roughgarden, 2009; Beyhaghi et al., 2021). However, these approaches rely on full knowledge of bidders’ valuation distributions, which is often unavailable or unreliable in practice. Moreover, they become computationally intractable in multi-unit auctions, as discussed earlier. In practice, historical bidding data is crucial for bidders to calibrate their strategies. For instance, Kasberger and Schlag (2024) note that none of the auction consultants surveyed could offer viable solutions without access to sufficient real-world data. These challenges, along with the increasing availability of auction data, have led to a growing interest in data-driven approaches. Beyond learning optimal bidding strategies in a repeated setting, as discussed in the previous section, this line of work includes data-driven reserve price optimization (Roughgarden and Wang, 2019; Derakhshan et al., 2022, 2021; Golrezaei et al., 2021a; Feng et al., 2021), boost value optimization in second-price auctions (Golrezaei et al., 2021b) and deep learning methods for auction design (Feng et al., 2018; Dütting et al., 2024; Wang et al., 2024, 2025; Liu et al., 2025).